# Dejavu: Towards Experience Feedback Learning for Embodied Intelligence

### v2

@@ B.2 Reward Shaping and SAC Training Details
This subsection makes explicit the implementation behind the similarity terms in Eq. 13 and the Soft Actor–Critic (SAC) [Haarnoja2018SAC] training procedure used for EFN.
@@
For reward shaping we instantiate $\mathrm{sim}(\cdot,\cdot)$ as a cosine based similarity on the per-patch vision tokens. Given two feature maps $X,Y\in\mathbb{R}^{256\times 4096}$ from the frozen VLA backbone, we first $\ell_{2}$-normalize each token along the channel dimension and form the cosine matrix
$S=\mathrm{token\_cosine\_matrix}(X,Y)\\ =\mathrm{norm}(X)\,\mathrm{norm}(Y)^{\top}\in[-1,1]^{256\times 256}.$ (23)
We then compute a Sinkhorn-regularized matching between the two token sets with a Gibbs kernel $K=\exp(S/\varepsilon_{\text{OT}})$, uniform row and column marginals, and $n_{\text{iter}}$ Sinkhorn iterations. The resulting transport plan $P\in\mathbb{R}^{256\times 256}$ satisfies approximately uniform marginals and we define the scalar score
$\mathrm{score}(X,Y)\;=\;\sum_{i,j}P_{ij}\,S_{ij},$ (24)
@@
$\mathrm{sim}(X,Y)\;=\;\tfrac{1}{2}\,\big(\mathrm{clip}(\mathrm{score}(X,Y),-1,1)+1\big).$ (25)
In all experiments we set $\varepsilon_{\text{OT}}=0.05$ and $n_{\text{iter}}=50$. No additional learned projection is applied before computing similarity; we rely on the backbone’s vision tokens and the Sinkhorn normalization to provide a stable semantic similarity in $[0,1]$. The same primitive $\mathrm{sim}(\cdot,\cdot)$ is used for $s_{t}^{\mathrm{next}},s_{t}^{\mathrm{cur}},s_{t}^{\mathrm{stay}}$ in Eq. 13.
@@
$s_{t}^{\mathrm{next}}$ $=\mathrm{sim}(\mathbf{F}_{t+1},\hat{\mathbf{F}}^{+}),$ (26)
@@
$a_{t}=s_{t}^{\mathrm{next}},p_{t}=s_{t}^{\mathrm{next}}-s_{t}^{\mathrm{cur}},m_{t}=1-s_{t}^{\mathrm{stay}},n_{t}=[\,\varepsilon-p_{t}\,]_{+},$ (27)
with tolerance $\varepsilon>0$ and $x]_{+}=\max(x,0)$. The shaped reward in Eq. 13 is then
$r_{t}=$ $w_{\text{abs}}\,a_{t}+w_{\text{prog}}\,[p_{t}]_{+}+w_{\text{mot}}\,m_{t}$ (28)
@@
$w_{\text{abs}}=0.1,w_{\text{prog}}=20,w_{\text{mot}}=0.3,w_{\text{lazy}}=20,$ (29)
$\varepsilon=5\times 10^{-3},\quad\lambda_{\text{time}}=0.01.$ (30)
These values give a reward distribution with non-trivial variance while keeping $r_{t}$ in a moderate range for SAC. In practice we tune the weights in the following way. We first fix $w_{\text{abs}}$ to a small constant that preserves the absolute similarity bonus without dominating other terms, then adjust $w_{\text{prog}}$ so that genuine progress toward the retrieved successor frame is reliably rewarded. If the agent tends to move but not approach the target view we slightly reduce $w_{\text{prog}}$ and increase $w_{\text{mot}}$. If the policy learns to idle at viewpoints that already match the target frame we increase $w_{\text{lazy}}$ (and, if necessary, $\varepsilon$) to make the anti-idling penalty stronger. The per-step time cost $\lambda_{\text{time}}$ is kept small and is only increased when trajectories become unnecessarily long. The residual $\ell_{2}$ penalty from the semantic reward in Sec  3.3 is disabled in our final experiments, that is we set $\lambda_{\text{res}}=0$ and let the tanh-squashed residual range control the magnitude of $\Delta\mathbf{a}_{t}$.
@@
We use a standard off-policy Soft Actor–Critic setup on top of the shaped reward in Eq. (28). The discount factor is $\gamma=0.98$. The SAC temperature $\alpha$ is adapted online to match a target entropy. We maintain $\log\alpha$ as a learnable scalar and update it with
$\mathcal{L}_{\alpha}=-\,\big(\log\alpha\big)\,\big(\log\pi_{\phi}(\Delta\mathbf{a}_{t}|\mathbf{c}_{t})+\mathcal{H}_{\text{target}}\big),$ (31)
@@
$\mathcal{H}_{\text{target}}=-\kappa\cdot(4\cdot 4096),\quad\kappa=0.25,$ (32)
@@
Both actor and critics are trained with Adam optimizers. We use learning rates
$\text{lr}_{\text{actor}}=3\times 10^{-4},\text{lr}_{\text{critic}}=3\times 10^{-4},\text{lr}_{\alpha}=3\times 10^{-4},$ (33)
@@
$\tau=0.005,\quad\bar{\theta}\leftarrow(1-\tau)\,\bar{\theta}+\tau\,\theta$ (34)
@@
Algorithm 1 summarizes the training loop that combines retrieval, reward shaping, and SAC. We emphasize that the frozen VLA backbone and experience bank are treated as environment-side components; gradients never flow into the backbone or the bank.
Algorithm 1 EFN training with reward shaping and SAC
@@
then construct $a_{t},p_{t},m_{t},n_{t}$ and reward $r_{t}$ according to Eq. (28).
@@
11:     Sample mini-batch from $\mathcal{D}$ and compute critic targets $y_{t}$ with soft Bellman backup using $Q_{\bar{\theta}_{i}}$ and $\alpha$ as in Sec  $3.$
@@
This procedure realizes the conceptual objective in Sec  3.3: EFN learns a residual policy that steers $\mathbf{F}_{t+1}$ toward the retrieved successor frame $\hat{\mathbf{F}}^{+}$ while discouraging degenerate idling and overly long trajectories.
@@ B.3 EFN Architecture and Implementation
EFN operates on the latent interface exposed by the frozen vision–language–action (VLA) backbone. At each non-blank step the backbone provides a set of visual patch embeddings of shape $(256,4096)$ and a small number of latent action tokens of shape $(4,4096)$. For the current state and the retrieved experience state, EFN receives both the visual embeddings and the latent tokens and predicts a residual in the same latent space, which is added to the backbone action tokens. All EFN parameters are trained while the backbone remains fixed, so only this residual branch is updated.
To keep the architecture lightweight yet expressive, we first compress the long visual sequence into a few learned latent vectors. This is implemented with a Multihead Attention Pooling (MAP) block inspired by Set Transformers and UniVLA [univla]. A bank of learned seed vectors plays the role of latent queries, while the per-patch visual embeddings act as keys and values. A custom attention module projects seeds and visual inputs into multihead query, key and value tensors, performs scaled dot-product attention, and projects the attended seeds back to the embedding dimension. Each MAP block is wrapped with RMS normalization and a position-wise feed-forward network. The feed-forward network uses a SwishGLU unit that applies a linear projection to twice the hidden size, splits it into a value and a gate, modulates the value with a SiLU-activated gate, and maps back to the embedding dimension. This yields compact latent representations of the current and retrieved views instead of operating directly on all $256$ visual patches.
Table 4: Per-step latency of EFN variants. The full actor–critic configuration is used as reference.
@@
The EFN policy network uses two MAP blocks to obtain visual latents for the current and retrieved states, each with four latent vectors of dimension $1024$. The current and retrieved latent action tokens are projected into the same $1024$-dimensional space. For cross-conditioning, EFN concatenates the pooled current visual latents, pooled retrieved visual latents and retrieved latent tokens along the sequence dimension, which yields a context sequence of length $12$. The projected current latent tokens serve as queries. A multihead attention layer with layer-normalized queries, keys and values computes an attended representation of the four query tokens over this $12$-token context. A small feed-forward block with normalization and dropout refines these tokens, and a two-layer Transformer encoder with pre-normalization further processes the four-token sequence to capture interactions among the latent dimensions. A final LayerNorm and a three-layer MLP map each token back to the original $4096$-dimensional latent space, with a $\tanh$ activation and explicit clipping to keep the residual inside a controlled range. The output is a residual tensor of shape $(4,4096)$ that is added to the backbone latent action tokens. At deployment we run this policy branch in inference mode and feed the refined latent tokens to the frozen VLA head without any further gradient updates.
@@
To quantify the overhead introduced by EFN itself, we benchmark several variants of the residual module under the same latent interface. Table 4 reports per-step latency for the full actor–critic configuration and for lighter policy-only variants. Removing the critic already reduces latency by more than $40\%$, and a one-layer encoder achieves the lowest cost while preserving the interface, which confirms that EFN adds only a modest runtime overhead on top of the frozen backbone.
@@ B.4 Simulation Setup and Baseline Configurations
We primarily evaluate on the LIBERO long-horizon manipulation benchmark [Liberoneurips2023], which provides language-conditioned tabletop tasks in a simulated kitchen. Following the official protocol, we use the four standard suites: LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long. Each suite groups tasks that emphasize spatial reasoning, object-centric manipulation, goal-directed rearrangement, and extended multi-step interactions, respectively. As reported in Table 1, our EFN variants consistently achieve the highest average success rate across all four suites under the same backbone and interaction budget.
We additionally validate EFN on the CALVIN language-conditioned tabletop manipulation benchmark [calvin], which focuses on multi-task goal sequences and compositional generalization. CALVIN differs from LIBERO in that an episode may consist of a chain of language goals and is evaluated with Success@k metrics over these goal chains. Due to space limits in the main paper, all quantitative and qualitative results on CALVIN, together with a detailed description of its task splits and evaluation metrics, are deferred to the supplementary material in Sec. E (Supplementary Experimental Results).
@@
Across all experiments, EFN wraps a frozen vision–language–action backbone. For the LIBERO simulations, we use the officially released checkpoints and configurations of OpenVLA [openvla] and UniVLA [univla] without architectural modifications. For the quadruped setting, we employ the publicly released AgiBot GO-1 [go1] controller as our backbone and keep its perception and low-level control stack unchanged. Its performance on LIBERO is reproduced by strictly following the official README instructions (training script, hyperparameters, and evaluation protocol); we do not perform additional tuning specific to our method, and all baselines share the same backbone initialization.
@@
Unless otherwise stated, all reported numbers are averaged over 50 evaluation episodes per task and 3 random seeds (i.e., 150 rollouts per cell in Table 1). During evaluation, policies act deterministically by using the mean action of the underlying stochastic policy (SAC) and the deterministic retrieval rule of each method. An episode is marked as Success if the environment-specific success flag is triggered before the task horizon; otherwise it is counted as a failure. The Step metric counts the number of control steps from the first non-blank action until success or timeout, where we ignore the initial blank warm-up steps (e.g., simulator loading and robot settling) that occur before the policy starts issuing meaningful actions. For CALVIN, we follow the official evaluation scripts and report Success@k and average sequence length metrics as defined in the benchmark, and success@k follows the official CALVIN definition: a sequence is successful if the first k sub-goals are completed in order.
@@
We implement a non-parametric kNN-RAG baseline that controls the agent purely by retrieving and reusing past actions from the experience bank, without learning any residual policy. At each non-blank step, we encode the current observation and language instruction with the same frozen VLA backbone used by all other methods and obtain a joint visual–language key. This key is compared to all step-level keys stored in the experience bank (Sec. 3.2) using cosine similarity, and we select the single nearest neighbor in this embedding space. The action stored in that retrieved transition is then directly executed as the agent’s action at the current step, i.e., the controller simply “copies what worked before” for the most similar past state.
@@
We further implement a ResAct-style residual RL baseline that adapts the core ideas of ResAct [resact], namely observation-difference features and action residuals, to our VLA setting. To keep the comparison fair, this baseline shares the same frozen vision–language–action backbone and residual RL infrastructure as our other methods. At each time step $t$, we obtain a visual embedding $f_{\text{vis}}(o_{t})$ from the VLA’s visual tokens (e.g., pooled over patches), along with proprioceptive features $f_{\text{prop}}(x_{t})$ and an instruction or goal embedding $e(\ell)$. In addition, we cache the previous-step visual embedding $f_{\text{vis}}(o_{t-1})$ and executed action $a_{t-1}$, and form an observation-difference feature
$d_{t}=f_{\text{vis}}(o_{t})-f_{\text{vis}}(o_{t-1}),$ (35)
@@
$s_{t}^{\text{ResAct}}=[\,f_{\text{vis}}(o_{t}),\ d_{t},\ a_{t-1},\ f_{\text{prop}}(x_{t}),\ e(\ell)\,],$ (36)
@@
$a_{t}=a_{t}^{\text{base}}+a_{t}^{\text{res}},$ (37)

### v3

@@ B.2 Reward Shaping and SAC Training Details
This subsection details the implementation of the similarity terms in main-paper Eq. (13) and the Soft Actor–Critic (SAC) [14] training procedure used for EFN.
@@
For reward shaping, we instantiate $\mathrm{sim}(\cdot,\cdot)$ as a cosine-based similarity on the per-patch vision tokens. Given two feature maps $X,Y\in\mathbb{R}^{256\times 4096}$ from the frozen VLA backbone, we first $\ell_{2}$-normalize each token along the channel dimension and form the cosine matrix
$S=\mathrm{token\_cosine\_matrix}(X,Y)\\ =\mathrm{norm}(X)\,\mathrm{norm}(Y)^{\top}\in[-1,1]^{256\times 256}.$ (B.7)
We then compute a Sinkhorn-regularized matching [10] between the two token sets with a Gibbs kernel $K=\exp(S/\varepsilon_{\text{OT}})$, uniform row and column marginals, and $n_{\text{iter}}$ Sinkhorn iterations. The resulting transport plan $P\in\mathbb{R}^{256\times 256}$ satisfies approximately uniform marginals and we define the scalar score
$\mathrm{score}(X,Y)\;=\;\sum_{i,j}P_{ij}\,S_{ij},$ (B.8)
@@
$\mathrm{sim}(X,Y)\;=\;\tfrac{1}{2}\,\big(\mathrm{clip}(\mathrm{score}(X,Y),-1,1)+1\big).$ (B.9)
In all experiments, we set $\varepsilon_{\text{OT}}=0.05$ and $n_{\text{iter}}=50$. No additional learned projection is applied before computing similarity; we rely on the backbone’s vision tokens and Sinkhorn normalization to provide a stable semantic similarity in $[0,1]$. The same primitive $\mathrm{sim}(\cdot,\cdot)$ is used for $s_{t}^{\mathrm{next}},s_{t}^{\mathrm{cur}},s_{t}^{\mathrm{stay}}$ in main-paper Eq. (13).
@@
$s_{t}^{\mathrm{next}}$ $=\mathrm{sim}(\mathbf{F}_{t+1},\hat{\mathbf{F}}^{+}),$ (B.10)
@@
$a_{t}=s_{t}^{\mathrm{next}},p_{t}=s_{t}^{\mathrm{next}}-s_{t}^{\mathrm{cur}},m_{t}=1-s_{t}^{\mathrm{stay}},n_{t}=[\,\varepsilon-p_{t}\,]_{+},$ (B.11)
with tolerance $\varepsilon>0$ and $x]_{+}=\max(x,0)$. The shaped reward in main-paper Eq. (13) is then
$r_{t}=$ $w_{\text{abs}}\,a_{t}+w_{\text{prog}}\,[p_{t}]_{+}+w_{\text{mot}}\,m_{t}$ (B.12)
@@
$w_{\text{abs}}=0.1,w_{\text{prog}}=20,w_{\text{mot}}=0.3,w_{\text{lazy}}=20,$ (B.13)
$\varepsilon=5\times 10^{-3},\quad\lambda_{\text{time}}=0.01.$ (B.14)
These values give a reward distribution with non-trivial variance while keeping $r_{t}$ in a moderate range for SAC. In practice, we tune the weights as follows. We first fix $w_{\text{abs}}$ to a small constant that preserves the absolute similarity bonus without dominating the other terms, then adjust $w_{\text{prog}}$ so that genuine progress toward the retrieved successor frame is reliably rewarded. If the agent tends to move without approaching the target view, we slightly reduce $w_{\text{prog}}$ and increase $w_{\text{mot}}$. If the policy learns to idle at viewpoints that already match the target frame, we increase $w_{\text{lazy}}$ (and, if necessary, $\varepsilon$) to strengthen the anti-idling penalty. The per-step time cost $\lambda_{\text{time}}$ is kept small and is only increased when trajectories become unnecessarily long. The residual $\ell_{2}$ penalty from the semantic reward in main-paper Sec. 3.3 is disabled in our final experiments, that is, we set $\lambda_{\text{res}}=0$ and let the tanh-squashed residual range control the magnitude of $\Delta\mathbf{a}_{t}$.
@@
We use a standard off-policy Soft Actor–Critic setup on top of the shaped reward in Eq. (B.12). The discount factor is $\gamma=0.98$. The SAC temperature $\alpha$ is adapted online to match a target entropy. We maintain $\log\alpha$ as a learnable scalar and update it with
$\mathcal{L}_{\alpha}=-\,\big(\log\alpha\big)\,\big(\log\pi_{\phi}(\Delta\mathbf{a}_{t}|\mathbf{c}_{t})+\mathcal{H}_{\text{target}}\big),$ (B.15)
@@
$\mathcal{H}_{\text{target}}=-\kappa\cdot(4\cdot 4096),\quad\kappa=0.25,$ (B.16)
@@
Both actor and critics are trained with Adam optimizers [19]. We use learning rates
$\text{lr}_{\text{actor}}=3\times 10^{-4},\text{lr}_{\text{critic}}=3\times 10^{-4},\text{lr}_{\alpha}=3\times 10^{-4},$ (B.17)
@@
$\tau=0.005,\quad\bar{\theta}\leftarrow(1-\tau)\,\bar{\theta}+\tau\,\theta$ (B.18)
@@
Algorithm B.1 summarizes the training loop that combines retrieval, reward shaping, and SAC. We emphasize that the frozen VLA backbone and experience bank are treated as environment-side components; gradients never flow into the backbone or the bank.
Algorithm B.1 EFN training with reward shaping and SAC
@@
then construct $a_{t},p_{t},m_{t},n_{t}$ and reward $r_{t}$ according to Eq. (B.12).
@@
11:     Sample a mini-batch from $\mathcal{D}$ and compute critic targets $y_{t}$ with the soft Bellman backup using $Q_{\bar{\theta}_{i}}$ and $\alpha$ as in main-paper Sec. $3.$
@@
This procedure realizes the conceptual objective in main-paper Sec. 3.3: EFN learns a residual policy that steers $\mathbf{F}_{t+1}$ toward the retrieved successor frame $\hat{\mathbf{F}}^{+}$ while discouraging degenerate idling and overly long trajectories.
@@ B.3 EFN Architecture and Implementation
EFN operates on the latent interface exposed by the frozen vision–language–action (VLA) backbone. At each non-blank step, the backbone provides a set of visual patch embeddings of shape $(256,4096)$ and a small number of latent action tokens of shape $(4,4096)$. For the current state and the retrieved experience state, EFN receives both the visual embeddings and the latent tokens and predicts a residual in the same latent space, which is added to the backbone action tokens. All EFN parameters are trained while the backbone remains fixed, so only this residual branch is updated.
To keep the architecture lightweight yet expressive, we first compress the long visual sequence into a few learned latent vectors. This is implemented with a Multihead Attention Pooling (MAP) block inspired by Set Transformers and UniVLA [8]. A bank of learned seed vectors plays the role of latent queries, while the per-patch visual embeddings act as keys and values. A custom attention module projects seeds and visual inputs into multihead query, key and value tensors, performs scaled dot-product attention, and projects the attended seeds back to the embedding dimension. Each MAP block is wrapped with RMS normalization and a position-wise feed-forward network. The feed-forward network uses a SwishGLU unit that applies a linear projection to twice the hidden size, splits it into a value and a gate, modulates the value with a SiLU-activated gate, and maps back to the embedding dimension. This yields compact latent representations of the current and retrieved views instead of operating directly on all $256$ visual patches.
Table B.1: Per-step latency of EFN variants. The full actor–critic configuration is used as reference.
@@
The EFN policy network uses two MAP blocks to obtain visual latents for the current and retrieved states, each with four latent vectors of dimension $1024$. The current and retrieved latent action tokens are projected into the same $1024$-dimensional space. For cross-conditioning, EFN concatenates the pooled current visual latents, pooled retrieved visual latents, and retrieved latent tokens along the sequence dimension, yielding a context sequence of length $12$. The projected current latent tokens serve as queries. A multihead attention layer with layer-normalized queries, keys, and values computes an attended representation of the four query tokens over this $12$-token context. A small feed-forward block with normalization and dropout refines these tokens, and a two-layer Transformer encoder with pre-normalization further processes the four-token sequence to capture interactions among the latent dimensions. A final LayerNorm and a three-layer MLP map each token back to the original $4096$-dimensional latent space, with a $\tanh$ activation and explicit clipping to keep the residual inside a controlled range. The output is a residual tensor of shape $(4,4096)$ that is added to the backbone latent action tokens. At deployment, we run this policy branch in inference mode and feed the refined latent tokens to the frozen VLA head without any further gradient updates.
@@
To quantify the overhead introduced by EFN itself, we benchmark several variants of the residual module under the same latent interface. Table B.1 reports per-step latency for the full actor–critic configuration and for lighter policy-only variants. Removing the critic already reduces latency by more than $40\%$, and a one-layer encoder achieves the lowest cost while preserving the interface, which confirms that EFN adds only a modest runtime overhead on top of the frozen backbone.
@@ B.4 Simulation Setup and Baseline Configurations
We primarily evaluate on the LIBERO long-horizon manipulation benchmark [25], which provides language-conditioned tabletop tasks in a simulated kitchen. Following the official protocol, we use the four standard suites: LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long. Each suite groups tasks that emphasize spatial reasoning, object-centric manipulation, goal-directed rearrangement, and extended multi-step interactions, respectively. As reported in Table 1 in the main paper, our EFN variants consistently achieve the highest average success rate across all four suites under the same backbone and interaction budget.
@@
Across all experiments, EFN wraps a frozen vision–language–action backbone. For the LIBERO simulations, we use the officially released checkpoints and configurations of OpenVLA [18] and UniVLA [8] without architectural modifications. For the real-world experiments, we employ the publicly released AgiBot GO-1 [7] controller as our backbone and keep its perception and low-level control stack unchanged. Its performance on LIBERO is reproduced by strictly following the official README instructions (training script, hyperparameters, and evaluation protocol); we do not perform additional tuning specific to our method, and all baselines share the same backbone initialization.
@@
Unless otherwise stated, all reported numbers are averaged over 50 evaluation episodes per task and 3 random seeds (i.e., 150 rollouts per cell in Table 1 in the main paper). During evaluation, policies act deterministically by using the mean action of the underlying stochastic policy (SAC) and the deterministic retrieval rule of each method. An episode is marked as Success if the environment-specific success flag is triggered before the task horizon; otherwise it is counted as a failure. The Step metric counts the number of control steps from the first non-blank action until success or timeout, where we ignore the initial blank warm-up steps (e.g., simulator loading and robot settling) that occur before the policy starts issuing meaningful actions.
@@
We implement a non-parametric kNN-RAG baseline that controls the agent purely by retrieving and reusing past actions from the experience bank, without learning any residual policy. At each non-blank step, we encode the current observation and language instruction with the same frozen VLA backbone used by all other methods and obtain a joint visual–language key. This key is compared to all step-level keys stored in the experience bank (main-paper Sec. 3.2) using cosine similarity, and we select the single nearest neighbor in this embedding space. The action stored in that retrieved transition is then directly executed as the agent’s action at the current step, i.e., the controller simply “copies what worked before” for the most similar past state.
@@
We further implement a ResAct-style residual RL baseline that adapts the core ideas of ResAct [29], namely observation-difference features and action residuals, to our VLA setting. To keep the comparison fair, this baseline shares the same frozen vision–language–action backbone and residual RL infrastructure as our other methods. At each time step $t$, we obtain a visual embedding $f_{\text{vis}}(o_{t})$ from the VLA’s visual tokens (e.g., pooled over patches), along with proprioceptive features $f_{\text{prop}}(x_{t})$ and an instruction or goal embedding $e(\ell)$. In addition, we cache the previous-step visual embedding $f_{\text{vis}}(o_{t-1})$ and executed action $a_{t-1}$, and form an observation-difference feature
$d_{t}=f_{\text{vis}}(o_{t})-f_{\text{vis}}(o_{t-1}),$ (B.19)
@@
$s_{t}^{\text{ResAct}}=[\,f_{\text{vis}}(o_{t}),\ d_{t},\ a_{t-1},\ f_{\text{prop}}(x_{t}),\ e(\ell)\,],$ (B.20)
@@
$a_{t}=a_{t}^{\text{base}}+a_{t}^{\text{res}},$ (B.21)